OpenStudy (anonymous):

Point E, F, G and H lie on a line that order. Ig EG=5/3EF and HF=5FG, then what is EF/HG?

OpenStudy (anonymous):

T_T Somebody please help me... This kind of problem had showed up several time on my tests in class and I get them wrong every time.. I would really appreciate it.

OpenStudy (anonymous):

wait can you rewrite the values of each of the line segments seperately, I'm a bit confused

OpenStudy (anonymous):

of course!\[EG=\frac{ 5 }{ 3 }EF\]\[HF= 5FG\] What is \[\frac{ EF }{ HG}\]

OpenStudy (anonymous):

Using the first equation, solve for EF. EG*(3/5)=(3/5)*(5/3)*EF EG*(3/5)=EF To find HG we can do HF-GF So HF = 5FG , FG = (1/5)HF 5FG-(1/5)(5FG)=5FG-FG = 4FG Now, rewrite.. EF/HG = (EG*(3/5))/5FG

OpenStudy (anonymous):

Thank you!!! :D

OpenStudy (anonymous):

No prob!! :D